OpenStudy (twalt13):

can anyone help me solve this problem? Give an example of a rational function that has a horizontal asymptote at y=1 and a vertical asymptote at x=4.

OpenStudy (tkhunny):

Getting a horizontal asymptote other than y = 0 requires the degree of the numerator to match the degree of the denominator. Examples: \[y = 1\] \[y = \frac{x-1}{x-2}\] Vertical Asymptote at x = 4 requires the denominator to be zero for x = 4. Example: \[y = \frac{x-1}{x-4}\]